OpenStudy (anonymous):

A standard deck of 52 cards has four 13-card suits: diamonds, hearts, clubs, and spades. The diamonds and hearts are red, and the clubs and spades are black. Each 13-card suit contains cards numbered from 2 to 10, a jack, a queen, a king, and an ace. The jack, queen, and king are called face cards. An experiment consists of randomly drawing a card from a standard deck. What is the probability that the card drawn is an ace?

OpenStudy (aaliyahfairgood):

4 out of 52 because there are 4 aces in a deck of cards ratio is 4:52 simplified ratio is 2:26

OpenStudy (anonymous):

~Since there are 4 aces in a standard deck and 52 cards in total : P(ace) = 4/52 = 1/13

OpenStudy (anonymous):

1/13 for ace and other numbers too.

OpenStudy (aaliyahfairgood):

That is simpler cool

OpenStudy (anonymous):

1/13 I guess!!!!!